### v1

# Depth-Breadth Synergy in RLVR: Unlocking LLM Reasoning Gains with Adaptive Exploration

Abstract
Reinforcement Learning with Verifiable Reward (RLVR) has emerged as a powerful paradigm for unlocking reasoning capabilities in large language models, yet its full potential is hindered by two under-explored dimensions: Depth—the hardest problem a model can sample; Breadth—the number of instances consumed in a single iteration. We dissect the popular GRPO algorithm and reveal a systematic bias: the cumulative-advantage disproportionately weights samples with medium accuracy, while down-weighting the low-accuracy instances that are crucial for pushing reasoning boundaries. To rectify the depth neglect, we introduce Difficulty Adaptive Rollout Sampling (DARS), which re-weights hard problems through targeted multi-stage rollouts, thereby increasing the number of positive rollouts for hard problems. Empirically, naively enlarging rollout size only accelerates convergence and even hurts Pass@K. Our DARS, in contrast, delivers consistent Pass@K gains without extra inference cost at convergence. Just as we adaptively expanded the depth of exploration, we now ask whether aggressively scaling the breadth of training data can further amplify reasoning gains. To this end, we intensely scale batch size and replace PPO’s mini-batch iterations with full-batch updates over multiple epochs. Increasing breadth significantly enhances Pass@1 performance. Large-breadth training sustains high token-level entropy, indicating continued exploration and reduced gradient noise. We further present DARS-B, which augments DARS with large breadth, and demonstrate simultaneous gains in Pass@K and Pass@1. The results confirm that breadth and adaptive exploration across depth operate as orthogonal dimensions in RLVR, which are key to unleashing the reasoning power of RLVR.
Figure 1: Training dynamics of Pass@1 and Pass@K performance. We show that our DARS significantly improves Pass@K performance and is complementary to large breadth scaling to further improve Pass@1 performance.
1 Introduction
The emergence of reasoning-centric Large Language Models (LLMs) exemplified by OpenAI-o1 (Jaech et al., 2024), DeepSeek-R1 (Guo et al., 2025), and Kimi-1.5 (Team et al., 2025), has pushed the frontier of LLM capability, especially for demanding tasks in complex reasoning such as mathematics and programming. Unlike conventional instruction tuning that relies on human-labeled data or RLHF pipelines that demand an auxiliary, well-trained reward model (Ouyang et al., 2022a; Achiam et al., 2023; Grattafiori et al., 2024), this leap is driven by large-scale Reinforcement Learning with Verifiable Rewards (RLVR; Guo et al. 2025; Zeng et al. 2025) for which correctness can be automatically and deterministically checked. The rewards of RLVR are granted solely when a model’s output matches the ground-truth mathematical answer or passes all unit tests for code, allowing scalable verification without manual labeling. RLVR is now regarded as a promising path toward continuously self-evolving LLMs, potentially bringing us closer to more powerful intelligence.
However, existing RLVR frameworks inadequately address the interplay between exploration depth (difficulty scaling) and breadth (iteration instance quantity scaling), which leads to insufficient performance gain for both Pass@1 and Pass@K. In this paper, we conduct a systematic analysis of two under-exploited dimensions in RLVR: Depth and Breadth.
For the dimension of depth, our investigation reveals that existing methods of GRPO (Shao et al., 2024) and its variants (Yu et al., 2025; Liu et al., 2025b), while adept at estimating the advantage of a single rollout, are undermined by a distorted cumulative advantage at the group level. This distortion disproportionately allocates attention to instances of medium difficulty, neglecting high-difficulty instances indispensable for complex reasoning, as illustrated in Figure 2. This bias fundamentally limits depth, the hardest problems a model can learn to solve, and constrains Pass@K performance. To counteract this depth neglect, we propose Difficulty-Adaptive Rollout Sampling (DARS). DARS performs a lightweight first-stage rollout to estimate per-problem accuracies, then allocates additional compute via targeted multi-stage rollouts to low-accuracy problems. By expanded sampling on hard problems, DARS re-weights the cumulative advantage, making it easier for LLMs to learn ‘deep’ samples and improving Pass@K performance.
We further identify breadth as the instance quantity consumed in a single iteration. We observe that breadth has a significant impact on the LLM’s performance and continuous exploration capability, as shown in Figure 5. We significantly increase the training batch size and replace PPO-minibatch updates with full-batch updates for multiple PPO-epochs. This seemingly simple change dramatically improves Pass@1 and sustains high token-level entropy throughout training, suggesting that breadth acts as implicit entropy regularization that delays premature convergence. Importantly, the gains from breadth are complementary to those from depth: we present DARS-B that combines our DARS with large-breadth training, producing simultaneous boosts in both Pass@K and Pass@1.
Our contributions can be summarized as follows:
  • We conduct a systematic analysis on depth and breadth in RLVR, and uncover the depth bias in GRPO: cumulative advantage silently underweights low-accuracy, high-difficulty samples, capping Pass@K performance.
  • We introduce Difficulty-Adaptive Rollout Sampling (DARS), which reallocates compute from medium difficulty problems to the hardest problems via multi-stage rollout sampling. DARS re-weights the cumulative advantage distribution and quantitatively expands the sparse reward signals for difficult problems. In practice, DARS significantly improves Pass@K over multiple benchmarks.
  • We further illustrate that large breadth in RLVR training matters for the Pass@1 performance. Moreover, by combining DARS with large breadth training, we reveal the complementarity of Depth and Breadth in RLVR and acquire simultaneous boosts in both Pass@K and Pass@1 performance.
2 Understanding RLVR from Depth and Breadth
In this section, we identify the two dimensions of Depth and Breadth and examine their impact on the Pass@1 and Pass@K metrics during the RLVR process. For training details and specific experimental settings, please refer to Section 4.1.
Figure 2: Statistical results of cumulative advantage. Both advantage calculation methods underestimate high-difficulty problems. $n$ denotes group size.
2.1 Depth: The Hardest Problem Sampled in RLVR
We first identify Depth as the hardest problem that can be correctly answered in the RLVR training process. In the GRPO training process, groups whose entire rollouts yield incorrect answers suffer from gradient vanishing. Hence, sampling high-difficulty questions with correct reasoning paths is crucial for LLM training. In this subsection, we first show that merely increasing rollout size does not consistently yield significant gains in Pass@K performance, and sometimes can even be harmful. We then quantify GRPO’s cumulative advantage and highlight its under-weighting of high-difficulty samples.
Figure 3: Training dynamics of Pass@1 and Pass@K performance of Qwen2.5-Math-1.5b and Qwen2.5-Math-7b with different rollout size.
Naive Scaling of Rollout Size Benefits Pass@1, But Not Necessarily Pass@K.
We present the training dynamics of Pass@1 and Pass@K performance during the RLVR training process in Figure 3. Enlarging the rollout size allows the sampling of correct solutions to hard problems during training. We originally assumed this would benefit Pass@K performance; however, experimental results show that this is not always the case. We find that Qwen2.5-Math-7b can significantly benefit from an increased rollout size, whereas for Qwen2.5-Math-1.5b, naively scaling rollout size can even harm Pass@K performance.
Cumulative Advantage Bias in GRPO Variants hinders the improvement of Pass@K.
In the GRPO (Shao et al., 2024) framework, the advantage estimation is derived by normalizing binary rewards:
$\hat{A}_{i}^{std}=\frac{r_{i}-u}{\sigma},~~~~~\hat{A}_{i}^{nostd}=r_{i}-u,$ (1)
where $r_{i}$ is the binary reward of $i_{th}$ rollout, $u$ is the mean value of the group rewards $u=\text{mean}(\{R_{i}\}_{i=1}^{G})$ and $\sigma$ is the standard deviation of the group rewards $\sigma=\text{std}(\{R_{i}\}_{i=1}^{G})$. In the case of binary rewards, $u$ also represents the accuracy of LLM rollouts. In Dr. GRPO (Liu et al., 2025b), the standard deviation term in the advantage is removed.
For a group $G$ with rollout size $N$, we define the cumulative advantage of a group as the sum of the absolute values of sample advantages: $\mathcal{A}_{\text{group}}\;=\;\sum_{i=1}^{G}\,|\hat{A}_{i}|$. Specifically,
$\mathcal{A}_{\text{group}}^{std}=2N\sqrt{u(1-u)},~~~\mathcal{A}_{\text{group}}^{nostd}=2Nu(1-u),$ (2)
Intuitively, cumulative advantage reflects how much the algorithm weights each sample. The cumulative advantage functional curve of GRPO style advantage and Dr. GRPO style advantage are plotted in Figure 2. As shown in the figure, group-based advantage computation funnels its weight toward problems of medium difficulty while largely overlooking those that are highly difficult. This bias limits the Pass@K performance of RLVR. To address this issue, we propose Difficulty Adaptive Rollout Sampling (DARS) to balance the cumulative advantage of hard samples, as detailed in Section 3.1.
Figure 4: Training dynamics of Pass@1 and Pass@K performance of Qwen2.5-Math-1.5b and Qwen2.5-Math-7b with different batch size.
2.2 Breadth: Iteration Instance Quantity in RLVR
We identify Breadth as the quantity of instances used in each iteration of the RLVR process. In this subsection, we discuss the Pass@1 performance benefits of simply increasing the batch size for the RLVR process. In section 4.3, we further discussed the impact of replacing the PPO mini batch with the full batch update for multiple PPO epochs.
Breadth Matters for Pass@1 Performance.
Most studies (Liu et al., 2025b; a; Yan et al., 2025; Fu et al., 2025) conventionally set the batch size to 128. In this subsection, we drastically increase the batch size to 3072 and plot the training dynamics of Pass@1 and Pass@32 performance in Figure 4. Naively increasing the batch size brings a Pass@1 improvement for all models, yet it harms the Pass@128 performance of Qwen2.5-Math-1.5b. We consider that increasing the quantity of instances used in each iteration makes the gradient direction more accurate and reduces the impact of noise, thereby improving Pass@1 performance.
Figure 5: Training dynamics of Pass@1 performance and token entropy for Qwen2.5-Math-1.5b and Qwen2.5-Math-7b.
Breadth Sustains Entropy for Model Exploration.
We further analyze the relationship between Pass@1 and token entropy during the training process, as shown in Figure 5. Higher token entropy signifies stronger exploration capabilities in LLMs. Increasing training breadth allows an LLM to achieve higher entropy at the same Pass@1 accuracy. We consider that large breadth serves as an entropy regularization mechanism, preventing premature convergence and boosting Pass@1 while sustaining entropy.
3 Methodology
In Section 2, we analyze the bias inherent in group-based advantage computation. To solve this issue, we introduce Difficulty Adaptive Rollout Sampling (DARS), which rebalances the cumulative advantage via multi-stage sampling. By further synergizing the depth and breadth training dimensions, we propose DARS-B, which improves both Pass@1 and Pass@K.
3.1 Difficulty Adaptive Rollout Sampling (DARS)
As illustated in Figure 6, given a data batch $\mathcal{B}=\{q_{j}\}_{j=1}^{M}$ of reasoning questions, DARS operates in two phases: (i) pre-rollout difficulty estimation step that assigns to each question $q_{j}$ a scalar difficulty score $x_{j}\in[0,1]$; and (ii) multi-stage rollout re-balancing step that dynamically decides how many additional trajectories $\Delta n_{j}$ shall be allocated to $q_{j}$ so that the cumulative advantage for low-accuracy problems is up-weighted.
To unify notation across the std-based and non-std advantage formulations, let
$\mathcal{S}(\hat{a}_{j})=\begin{cases}2\sqrt{\hat{a}_{j}(1-\hat{a}_{j})},&\text{for std-based advantage},\\[4.0pt] 2\hat{a}_{j}(1-\hat{a}_{j}),&\text{for non-std advantage}.\end{cases}$ (3)
Phase 1: Pre-Rollout Difficulty Estimation.
For every $q_{j}$, we draw a light first-stage rollout consisting of $k_{0}$ independent trajectories $\{\tau_{j}^{(i)}\}_{i=1}^{k_{0}}$. Let the per-trajectory reward be binary, $r_{j}^{(i)}\in\{0,1\}$. We define the empirical accuracy
$\hat{a}_{j}=\frac{1}{k_{0}}\sum_{i=1}^{k_{0}}r_{j}^{(i)}.$ (4)
The difficulty score is then set to the complementary accuracy $x_{j}=1-\hat{a}_{j}$, so that $x_{j}\approx 1$ for the hardest problems and $x_{j}\approx 0$ for the easiest ones.
Phase 2: Multi-Stage Rollout Re-Balancing.
Let $\mathcal{A}_{\text{group}}^{N}(u)$ denote the cumulative advantage under GRPO for a group whose average accuracy is $u$ with rollout size $N$. We aim to reallocate $\Delta N$ additional trajectories across the mini-batch so that the effective cumulative advantage for each question becomes an increasing function of its difficulty. To control the computational cost of rollout sampling, we set the maximum sampling upper limit $N^{\mathrm{max}}$. To this end, we design two re-balancing schedules.
Schedule 1: Equal-Treatment (ET). For every question $q_{j}$ we enforce the rebalanced cumulative advantage as:
$\mathcal{A}_{\text{group}}^{ET}(q_{j})=\mathcal{A}_{\text{group}}^{N}(0.5).$ (5)
We raise the cumulative advantage of all difficulty problems ($\hat{a}_{j}<0.5$) to the level achieved by a medium-difficulty problem with accuracy $\hat{a}_{j}=0.5$. The required extra trajectories are
$\Delta n_{j}^{\text{ET}}=\mathrm{min}(\left\lceil\frac{\mathcal{A}_{\text{group}}^{N}(0.5)-\mathcal{A}_{\text{group}}^{N}(\hat{a}_{j})}{\mathcal{S}(\hat{a}_{j})}\right\rceil,N^{\mathrm{max}}).$ (6)
Schedule 2: Hardness-Weighted (HW). We now impose a monotonically increasing re-weighting that allocates more rollouts to lower-accuracy problems:
$\mathcal{A}_{\text{group}}^{HW}(q_{j})=2(1-x_{j})\mathcal{A}_{\text{group}}^{N}(0.5).$ (7)
This yields
$\Delta n_{j}^{\text{HW}}=\mathrm{min}(\left\lceil\frac{2x_{j}\cdot\mathcal{A}_{\text{group}}^{N}(0.5)-\mathcal{A}_{\text{group}}^{N}(\hat{a}_{j})}{\mathcal{S}(\hat{a}_{j})}\right\rceil,N^{\mathrm{max}}).$ (8)
Figure 6: The overall training framework of our Difficulty Adaptive Rollout Sampling (DARS) with breadth scaling. Our DARS consists of 2 phases: 1) a pre-rollout stage to evaluate the difficulty of the given question, and 2) a re-balancing rollout stage to adjust the cumulative advantage. For breadth scaling, we replace ppo minibatch as fullbatch with multiple ppo epochs.
3.2 Depth Synergy with Breadth Scaling
Our analysis in Section 2.2 empirically confirms the substantial Pass@1 improvements from large-breadth training. While DARS primarily optimizes training depth via multi-stage rollout rebalance, its dynamic batch-size adjustments preclude standard PPO-style mini-batch updates. To resolve this architectural constraint while leveraging breadth benefits, we replace PPO’s mini-batch updates with full-batch gradient descent across multiple PPO epochs, as illustrated in Figure 6. This modification ensures compatibility with DARS’s dynamic allocation while maximizing effective training breadth per optimization step. We term this integrated approach DARS-B, unifying depth-adaptive sampling with breadth maximization.
Full-batch training offers two key advantages: (1) elimination of mini-batch gradient noise, and (2) sustained token-level exploration, acting as implicit regularization against premature convergence. The resulting framework demonstrates complementary gains—DARS improves Pass@K through depth optimization, while large-breadth training enhances Pass@1, highlighting their synergistic roles in RLVR optimization.
3.3 Training Target
We adopt the clipped objective of GRPO without the KL penalty term. Following Dr. GRPO, we likewise remove the response length handling from the GRPO target. Specifically, for a problem $q$ sampled in training data $\mathcal{D}$, the training target is formalized as:
$\mathcal{J}$ $(\theta)=\mathbb{E}_{(q\sim\mathcal{D},\{o_{i}\}_{i=1}^{\mathcal{G}}\sim\pi_{\theta_{\text{old}}}(q)}$ (9)
$\Bigg{[}\frac{1}{G}\sum_{i=1}^{G}\sum_{t=1}^{|o_{i}|}\Bigg{(}\min\Big{(}r_{i,t}(\theta)\hat{A}_{i,t},\ \text{clip}\Big{(}r_{i,t}(\theta),1-\varepsilon,1+\varepsilon\Big{)}\hat{A}_{i,t}\Big{)})\Bigg{)}\Bigg{]},$
where
$r_{i,t}(\theta)=\frac{\pi_{\theta}(o_{i,t}\mid q,o_{i,<t})}{\pi_{\theta_{\text{old}}}(o_{i,t}\mid q,o_{i,<t})}.$ (10)
The token advantage $\hat{A}_{i,t}$ is computed using Equation 1.
4 Experiments
4.1 Setup
Evaluation and Training.
We evaluate the RLVR process using 5 widely used mathematical reasoning benchmarks: MATH-500 (Lightman et al., 2023), OlympiadBench (He et al., 2024), MinvervaMath (Lewkowycz et al., 2022), AIME24, and AMC23. Math-500 contains 500 items screened out from the original MATH test split. The subset consists of 500 representative problems, and the evaluation produces similar results to the full-set evaluation. We combine all of the evaluation benchmarks to report Pass@K performance. The training data we used in this work is a subset of OpenR1-Math-220k (Hugging Face, 2025). Following LUFFY (Yan et al., 2025), we use the default subset and filter out generations that are longer than 8192 tokens and those that are verified wrong by Math-Verify  [^1] https://github.com/huggingface/Math-Verify, resulting in 45k question-solution pairs.
Parameters and Metrics.
Currently, our experiments are conducted with Qwen2.5-Math series language models (Yang et al., 2024). We set the temperature to 1.0 for both training and evaluation procedure. In this paper, we mainly use two metrics, Pass@1 and Pass@K. To acquire Pass@K results, we sample 128 candidate responses for each question during the evaluation process; the calculation of Pass@1 is derived from Avg@128. Both the training and evaluation processes are scored using Math-Verify. The learning rate is 1e-6 for depth training methods, and 5e-6 for large breadth training. We do not use the reference model and KL loss. For fair comparison, we uniformly set the PPO mini step to 2 for all experiments. By default, the maximum prompt length is 1024, and the maximum response length is 3072 for the Qwen2.5-Math series model.
Implementation Details.
Our training framework is derived from Verl (Sheng et al., 2024) pipeline, which is a flexible, high-performance reinforcement-learning framework built for training large language-model agents. With native PyTorch support and efficient distributed training, Verl lets researchers quickly prototype and scale RL algorithms like PPO on GPUs. All of our experiments are conducted on H200 GPUs. At present, the LLM of our experiment is the Qwen2.5-Math series. Due to the limitations of time and computation resources, we have only reported the 1.5b and 7b models. We plan to conduct experiments on more LLMs soon. More details are included in Appendix A.1.
Base Model + Method AIME24 MATH500 Olympiad AMC Minerva Avg@128 Pass@128
Qwen2.5-Math-1.5B 4.0 35.1 16.2 20.8 9.5 21.1 77.9
GRPO baseline 14.7 75.9 39.4 47.5 31.2 49.6 79.6
Depth-Naive 16.5 76.2 39.9 47.9 30.9 50.1 79.9
Breadth-Naive 18.5 77.6 41.7 49.8 31.9 51.5 79.2
DARS (ours) 15.8 76.0 40.9 47.2 30.0 50.0 81.2
DARS-B (ours) 18.6 79.4 42.9 50.6 31.7 52.5 80.8
Qwen2.5-Math-7B 11.6 52.3 19.7 35.2 15.3 30.1 82.1
GRPO baseline 26.8 82.2 44.3 57.2 35.7 55.3 81.4
Depth-Naive 28.0 83.8 46.4 59.0 37.3 57.0 80.3
Breadth-Naive 30.5 83.7 47.3 61.4 37.7 57.7 79.2
DARS (ours) 26.9 83.2 46.6 57.3 38.5 57.0 81.7
DARS-B (ours) 33.3 83.8 47.8 61.3 38.4 58.1 82.1
Table 1: Overall performance of Pass@1 (Avg@128) and Pass@128 performance.
4.2 Main Results
Baseline and Methods:
We compare the following methods: (1) GRPO-baseline: GRPO with rollout size 8 and batch size 128. (2) Depth-Naive: Simply increasing the rollout size to 32. (3) Breadth-Naive: Simply increasing the batch size to 3072. (4) DARS: Our algorithm introduced in Section 3.2, using batch size 128 and $N^{\max}=32$. (5) DARS-B: Our Depth-and-Breadth synergy algorithm introduced in Section 3.3, using batch size 3072 and $N^{\max}=32$. For all methods, the number of PPO mini-steps was uniformly set to 2.
Evaluation Protocol:
For each method, we select the checkpoint with the highest Pass@1 accuracy for reporting. Table 1 summarizes the Avg@128 performance on each benchmark, the overall Pass@1 across all test data, and the Pass@128 performance.
Analysis on experimental results:
  • Breadth scaling significantly benifit Pass@1 performance. Across all model scales and tasks, Breadth-Naive surpasses both GRPO baseline and Depth-Naive in average Pass@1 (Avg@128). For instance, Breadth-Naive improves over GRPO baseline by 1.9–3.7 points on AIME24, MATH500, and Olympiad benchmarks.
  • Synergy emerges from complementary strengths. The consistent margin of DARS-B over both Breadth-Naive and DARS validates our core hypothesis: depth and breadth are not competing resources; rather, their synergy is essential for unlocking LLM reasoning gains.
  • While Pass@1 reflects practical single-shot performance, Pass@128 measures the upper-bound capability under heavy sampling. DARS-B achieves the highest Pass@1 performance and matches the top Pass@128 scores. This indicates that our adaptive strategy not only improves the likelihood of finding a correct solution quickly but also raises the overall solution density within the search space.
Figure 7: Training dynamics of Pass@128 performance with different training steps of Qwen2.5-Math-1.5b and Qwen2.5-Math-7b.
4.3 Training Dynamics and Ablation Study
In this subsection, we further show more training dynamics to illustrate properties of existing RLVR methods and the superiority of our DARS and DARS-B.
Pass@128 performance surpasses the base model, peaks quickly, and then declines.
We conduct RLVR experiments with rollout size 8/32 to compare our DARS (with $N^{\mathrm{max}}=32$ for fair comparison), the training dynamics of Pass@128 performance during training is shown in Figure 7. Across all settings, Pass@128 surpasses the base model during training, but declines after peaking, indicating that over-training with RLVR harms Pass@128 performance. Throughout training, our DARS not only achieves the highest peak Pass@128 performance but also surpasses all other settings.
Figure 8: Training dynamics of Pass@32/Pass@128 and Pass@1 performance with different training steps of Qwen2.5-Math-1.5b and Qwen2.5-Math-7b.
Depth Training with DARS Improve Pass@K performance.
Because the Pass@K (K=32/128) metric is hard to improve monotonically—it even starts to drop after prolonged training—while Pass@1 remains comparatively stable and rarely collapses, we seek to boost Pass@K without degrading Pass@1. Figure 8 below plots Pass@K against Pass@1 under a variety of experimental settings. It shows that, at any fixed Pass@1 level, our DARS method delivers a consistently higher Pass@K than the other settings. It is worth noting that, unlike the naive approach of simply increasing the rollout size to 32, our DARS achieves significantly higher training efficiency by allocating more rollouts only to the hard problems.
Depth-Breadth are Complementary in RLVR.
We further illustrate that Depth and Breadth are two complementary dimensions in RLVR. We present DARS-B as illustrated in Section 3.2. As shown in Figure 9, we present the Pass@1Pass@K training-dynamics curves for the Breadth, Depth, and the two-dimensional synergy approach DARS-B. The farther the Pass@1Pass@K curve deviates outward, the more powerful the method. Our DARS-B curves lie on the outermost envelope: DARS-B not only achieves the best Pass@1, it simultaneously lift Pass@K. This demonstrates the complementary roles of Depth and Breadth in RLVR training. Moreover, we also select the best checkpoint of Pass@1 accuracy for each method, the performance comparison in each benchmark is shown in Table 1.
Figure 9: Complementary improve of Depth and Breadth Synergy for Pass@1 and Pass@K (K=32/128) performance.
5 Related Works
5.1 RL for LLM Reasoning
Reinforcement Learning (RL) is widely used in the post-training phase of LLMs. Previously, a reward model was employed to conduct preference comparisons among responses (Ouyang et al., 2022b). Direct Preference Optimization (Rafailov et al., 2023) is further proposed to simplify the training pipeline of RL for LLMs, which directly uses pair-wise preference data for model optimization. Recent studies have shown that RLVR can significantly improve the reasoning capabilities of LLMs in domains such as mathematics and programming. OpenAI’s o1 model (Jaech et al., 2024) represents a major milestone in LLM reasoning. Shortly thereafter, DeepSeek-R1 (Guo et al., 2025) demonstrated that RLVR can directly elicit a base model’s slow-thinking capability—a setting now termed zero-RL (Li et al., 2025). Subsequently, a series of Large Reasoning Models (LRM), including Kimi 1.5 (Team et al., 2025), Gemini-Think (DeepMind, 2024), and QwQ (Qwen, 2024) were successively introduced. These works have inspired a wave of follow-up research: Zeng et al. (2025) conducted extensive experimental exploration of zero-RL, while Luo et al. (2025) leveraged RLVR to further boost the performance of open-source models developed from DeepSeek-R1. Among these works, one widely adopted RLVR algorithm is GRPO (Shao et al., 2024), which builds upon PPO (Schulman et al., 2017) and attains substantial performance gains by sampling multiple responses to compute the group-relative advantage. GRPO has inspired a series of subsequent algorithms, including DAPO (Yu et al., 2025), VAPO (Yue et al., 2025b), and Dr. GRPO (Liu et al., 2025b). However, both GRPO and its variants tend to undervalue high-difficulty problems under cumulative advantage, which compromises the Pass@K performance of RLVR.
5.2 Analysis of RLVR
With the rapid advancement of RLVR and the proliferation of open-source LRMs, many studies have begun to analyze the RLVR pipeline and these open LRMs. Several studies (Liu et al., 2025a; Zhao et al., 2025; Shah et al., 2025) indicates that the self-reflect and self-critique behaviors observed after RLVR originates from the base model rather than the RL process. Dang et al. (2025) find that although the RLVR process benefits Pass@1, Pass@K may decline as training progresses. Subsequently, Yue et al. (2025a) through extensive experimental analysis, discovered that RLVR’s performance is significantly constrained by the base model; once training converges, it struggles to surpass the capability boundary of the base model. These studies have sparked widespread concern about the capability ceiling of RLVR, and consequently, the Pass@K metric has become a focal point for diagnosing and potentially transcending the intrinsic limits imposed by the base model. This paper analyzes and refines the RLVR pipeline from the dual perspectives of Pass@1 and Pass@K.
6 Conclusion
In this work, we identify and systematically analyze two critical yet under-explored dimensions in Reinforcement Learning with Verifiable Reward (RLVR): Depth (the hardest problems a model can learn) and Breadth (the number of instances processed per iteration). We reveal that existing GRPO-based methods suffer from a fundamental limitation—their cumulative advantage calculation disproportionately weights medium-difficulty problems while neglecting hard ones, creating a bottleneck for Pass@K performance. To address this issue, we propose Difficulty-Adaptive Rollout Sampling (DARS) to reallocate compute to the hard problems, boosting Pass@K with low extra inference cost. Furthermore, we demonstrate that large-breadth training acts as an orthogonal dimension for RLVR optimization. Scaling batch size and employing full-batch updates significantly boost Pass@1 performance while sustaining high token-level entropy. By unifying these dimensions in our DARS-B framework, we achieve simultaneous gains in both Pass@1 and Pass@K, confirming the synergistic relationship between depth and breadth in RLVR.
Appendix A Appendix
A.1 Setting Details
[…]